# Methodology, Data Curation, Formal analysis, Experimentation, Writing - Original Draft

### v1

@@ 1 Introduction
    We present a novel framework "XG-NID" that fuses flow-level and packet-level data modalities in network traffic, providing contextual, explainable, and actionable insights to enhance model interpretability and support informed decision-making.
@@
    We introduce an innovative method for fusing dual modalities of network traffic—flow and packet-level information—using heterogeneous graph structures for comprehensive analysis.
@@
    We develop a heterogeneous graph neural network (HGNN) based on graph-level classification for NIDS, enabling real-time inference by utilizing dual-modality data and incorporating the heterogeneous nature of network traffic.
@@
    We develop "GNN4ID," (Farrukh et al., 2024b) an open-source tool for extracting and transforming raw network traffic into our proposed heterogeneous graph structure, integrating flow and packet-level data.
  6. 5.
    We propose a new set of flow features based on temporal information to aid in producing better contextual and explainable inferences.
  7. 6.
    We present a methodology for providing detailed and human-readable explanations of the proposed HGNN model results by combining contextual information from LLMs.
  8. 7.
    We provide a quantitative comparative analysis of our proposed methodology with other baseline and comparable approaches to highlight its effectiveness.
@@ 2.2 Explainable Artificial Intelligence
Fig. 1.: Illustration of the proposed framework ”XG-NID”. The framework, depicted here, demonstrates its capability to better identify network attacks either based on flow or packet-level information with detailed contextual information and potential remedial actions. The framework comprises of six key components: Flow and Feature generator, Explainable Feature Extractor, Graph Generator, GNN Model, Integrated Gradient Explainer, and Generative Explainer.
@@ 3.1 Proposed Framework
The proposed framework ”XG-NID” presents a novel approach to network intrusion detection by integrating flow and packet-level information into a heterogeneous graph structure, enabling a more comprehensive analysis of network behavior. This fusion of multiple data modalities enhances the model’s ability to capture complex interactions within the network, leading to more accurate and effective intrusion detection. Moreover, the framework not only provides performant detection capabilities but also delivers detailed, human-readable explanations and potential remedial actions, ensuring that the insights generated are both actionable and understandable. The framework is composed of six key components, as illustrated in Fig. 2.2.
The system initiates by capturing raw network traffic data (network packets), which are then processed by the first component, the Flow and Feature Generator. This component is responsible for generating network flows and extracting relevant features from both the flow and packet levels. These generated flows are subsequently passed to the Explainable Feature Extractor, which derives new features based on the temporal information from previous flows.
@@
Once the features have been generated and enhanced, the data is transferred to the Graph Generator. This component transforms the flow and packet-level features into a heterogeneous graph structure, which serves as the foundation for the subsequent inference process. The transformed graph is then analyzed by the GNN Model, the fourth component, which performs the core task of intrusion detection by classifying the network traffic as either benign or respective attack class.
The output of the GNN Model is forwarded to both the Integrated Gradient Explainer and the Generative Explainer, components five and six, respectively. The Integrated Gradient Explainer identifies the most significant attributes of the graph structure that contributed to the model’s prediction, providing a localized explanation of the model’s output.
Leveraging these local explanations, the Generative Explainer generates specific prompts and formulates detailed, human-readable explanations of the results. Also, it offers potential remedial actions, ensuring that the system’s outputs are not only accurate but also actionable and understandable.
The following sub-sections provide the technical details of each component, offering a comprehensive understanding of how our framework is utilized for effective cyber threat detection and explanation.
@@ 3.1.6 Generative Explainer
Developing a structured prompt is particularly critical in zero-shot prompting approach because it allows the LLM to generate accurate and contextually relevant explanations without requiring additional training on specific examples. The effectiveness of zero-shot prompting relies heavily on the quality of the prompts. Therefore, developing well-structured prompts is essential to guide the LLM in producing meaningful outputs. The process of crafting these prompts is detailed in Algorithm 3.
@@
These parts—($P_{\text{init}}$, $P_{\text{part2}}$, and $P_{\text{align}}$)—are then combined to form a comprehensive query ($Q_{\text{flow}}$), which is then sent to the LLM to generate the flow-based response. If the GNN predicts a payload-specific attack, such as web-based or bruteforce attacks, additional processing is done for the payload data. In this processing, the payload importance vectors are normalized, and the top payloads are converted into a human-readable ASCII string. Consequently, a second query ($Q_{\text{payload}}$) is constructed to analyze the payload data. This query is prefixed with the phrase assigned to $P_{\text{payloadPrefix}}$: “Analyze whether this payload of network flow is malicious or not. Give reason concisely.” This query, along with the alignment section ($P_{\text{align}}$), is sent to the LLM to generate a corresponding response.
@@ 3.2.1 Dataset Preprocessing
The second task involved addressing the significant class imbalance present in the CIC-IoT2023 dataset. Given the uneven distribution of classes, a combination of undersampling and oversampling techniques was employed to create a balanced dataset. Initially, 20% of the data samples were set aside to form the test set, with majority classes capped at 4,000 samples for this purpose. The remaining data were then adjusted, either through undersampling the majority classes or oversampling the minority classes, to construct a training set of 20,000 samples per class. This approach ensured a more balanced distribution of samples across all classes, as illustrated in Table 4.
@@
Type Flows After Filter Test Set (20%) Train Set (80%)
@@
An important aspect of our preprocessing approach was the careful handling of attack subclasses. Since each main attack class in the CIC-IoT2023 dataset contains several subclasses, we ensured that the sampling process maintained proportional representation across these subclasses. This approach allowed us to preserve the diversity of attack types while achieving a more coherent and balanced dataset.
@@ 4 Results and Analysis
This section presents the results and analysis of our proposed framework, divided into two sub-headings: Performance Analysis and Explainability Analysis. The Performance Analysis subsection provides a comprehensive evaluation of the proposed approach, comparing its effectiveness against baseline models and state-of-the-art techniques. This comparison is aimed at demonstrating the superiority and robustness of our model in detecting network intrusions by fusing packet and flow level information together. Conversely, the Explainability Analysis subsection explores how our framework leverages contextual information to produce explainable and actionable insights, thereby enhancing the interpretability of the model’s predictions and facilitating informed decision-making by security operators.
It is important to note that all results presented in this section are based on multi-class classification, encompassing eight distinct categories from the CIC-IoT2023 dataset. These evaluations were conducted under consistent experimental settings, ensuring identical data distribution and processing across all evaluation.
@@ 4.1 Performance Analysis
Conducting an apple-to-apple comparison for the proposed framework presents certain challenges, as our work is the first to fuse dual modalities—packet-level and flow-level information of network traffic—through a heterogeneous GNN, as per the authors knowledge. Consequently, comparing our framework with other approaches that solely utilize either flow-level or packet-level information may not provide an accurate assessment of its effectiveness. Furthermore, comparing our proposed HGNN with other graph-based approaches is also problematic, as the majority of existing works in NIDS utilizing GNNs focus on edge or node-level classification. These methods primarily aim at analyzing network traffic rather than providing real-time inference, as highlighted in Section 2. Therefore, comparing our graph-level classification with such node and packet-level classification makes no sense.
Given that our approach emphasizes real-time inference with a focus on effectively fusing dual modalities of network traffic to improve detection of both flow-level and packet-level attacks, a direct comparison with existing methods is inherently complex. Therefore, to provide a comprehensive evaluation of our approach, we divided our performance analysis into two distinct parts.
@@ 4.1.1 Baseline Comparisons
In evaluating the baseline models, our objective is to offer a holistic view of the proposed framework when compared to approaches that solely rely on either flow-level or packet-level information. To provide deeper insights into the performance of the proposed framework, we also assessed the effectiveness of these baseline models on specific attack types—namely, payload-dependent attacks (such as brute force and web-based attacks) when relying on flow-level information, and payload-independent attacks when relying on packet-level information.
@@
Fig. 3.: Performance comparison between the proposed framework and several state-of-the-art models that incorporate dual modalities of network traffic. The proposed framework outperforms all other approaches, achieving a 97% F1-Score—a metric widely recognized for evaluating multi-class classification.
@@
AEIDS (Pratomo et al., 2018) 0.79 0.64 0.72 0.64
Packet2Vec (Goodman et al., 2020) 0.68 0.6 0.62 0.62
EsPADA (Vidal et al., 2020) 0.64 0.6 0.62 0.58
@@ 4.1.2 Dual-Modality State-of-the-Art
When compared with these state-of-the-art methods, our proposed framework demonstrates superior performance, as highlighted in Fig. 4.1.1. The inherent fusion of packet-level and flow-level information in our model allows for more accurate and efficient classification, setting a new benchmark in network intrusion detection. Our results clearly indicate that by integrating both modalities within a single framework, we can achieve better detection rates and more actionable insights than existing approaches, solidifying our framework’s position as a cutting-edge solution in the field.

### v2

@@ 1 Introduction
    We present XG-NID, a novel framework that fuses flow-level and packet-level data modalities using a heterogeneous graph structure, enabling real-time, context-aware network intrusion detection with enhanced explainability and actionable insights.
@@
    We develop a Heterogeneous Graph Neural Network (HGNN) model for graph-level classification in NIDS, designed to capture complex network interactions and support real-time inference, while accommodating the diverse nature of network traffic.
@@
    We introduce GNN4ID (Farrukh et al., 2024b), an open-source tool that transforms raw network traffic into the proposed heterogeneous graph format, streamlining the integration of flow and packet-level data for comprehensive analysis.
@@
    We propose new temporal flow features to improve contextual understanding and explainability in network intrusion detection, and leverage LLMs to generate detailed, human-readable explanations and potential mitigation strategies.
@@ 2.2 Explainable Artificial Intelligence
Fig. 1.: Illustration of the proposed framework ”XG-NID”. The framework, depicted here, demonstrates its capability to better identify network attacks either based on flow or packet-level information with detailed contextual information and potential remedial actions. The framework comprises of six key components: Flow and Feature generator, Explainable Feature Extractor, Graph Generator, HGNN Model, Integrated Gradient Explainer, and Generative Explainer.
@@ 3.1 Proposed Framework
The proposed framework, XG-NID, introduces a novel approach to network intrusion detection by integrating flow-level and packet-level information into a heterogeneous graph structure. This multi-modal data fusion empowers the model to capture the complex interactions within network traffic, enabling more accurate, robust, and context-aware intrusion detection. By incorporating both granular (packet-level) and aggregated (flow-level) data, XG-NID addresses the limitations of traditional NIDS and provides a more holistic view of network behavior. A key strength of XG-NID lies in its dual focus on high-performance detection and interpretability. Beyond achieving state-of-the-art accuracy, the framework prioritizes human-readable explanations and actionable insights to bridge the gap between technical detection outputs and decision-making by security analysts or non-expert users. This is achieved through the integration of a LLM, which enhances the explainability of detected threats, making complex cyber threat information accessible even to users without cybersecurity expertise.
The name XG-NID reflects these core principles: ’X’ stands for Explainable, emphasizing the framework’s focus on generating understandable insights; ’G’ denotes the use of a GNN for modeling complex network interactions; and ’NID’ represents the Network Intrusion Detection System, the primary application domain. The ’X’ particularly highlights the role of the LLM in providing contextualized explanations and remedial guidance based on the HGNN’s detection results.
The framework is designed as a modular pipeline comprising six key components, each playing a distinct yet interconnected role in the intrusion detection process (as illustrated in Fig. 2.2). The process begins with the Flow and Feature Generator, which captures raw network traffic and processes it to generate network flows while extracting both flow-level and packet-level features. This ensures that the system has a rich representation of network activity, essential for detecting both broad-scale and fine-grained attacks. These generated flows are subsequently passed to the Explainable Feature Extractor, which derives new temporal features based on historical flow data. By tracking patterns over time, this component enables the detection of time-based anomalies and strengthens the system’s ability to explain its decisions.
@@
Once the features have been generated and enhanced, the data is transferred to the Graph Generator, which transforms the flow and packet-level features into a heterogeneous graph structure. In this graph, nodes represent flows and packets, while edges capture relationships such as packet sequences and flow connections, providing a comprehensive foundation for the inference process. This transformed graph is then analyzed by the core detection engine of the framework, the HGNN Model. The HGNN processes the graph to classify network traffic as either benign or belonging to specific attack classes. Its ability to model both node-level and edge-level information allows it to capture the nuanced behaviors of complex cyber threats.
The output of the HGNN Model is forwarded to both the Integrated Gradient Explainer and the Generative Explainer. The Integrated Gradient Explainer applies Integrated Gradients to the HGNN’s predictions, identifying the most significant attributes of the graph structure that contributed to the model’s prediction. This step provides a transparent view into the inner workings of the model, offering localized explanations of its outputs. Building upon these feature attributions, the Generative Explainer uses an LLM to convert technical explanations into clear, human-readable insights. It also provides potential mitigation strategies, ensuring that the system’s outputs are not only accurate but also actionable and understandable, even for individuals without specialized security knowledge.
This structured flow—from raw data ingestion to actionable insights—ensures that XG-NID is both technically rigorous and practically viable. The modular design allows for flexible integration and future enhancements, while the dual-modality fusion and explainability focus set a new standard for modern NIDS. The following sub-sections provide a detailed technical breakdown of each component, highlighting their individual roles and the interdependencies that contribute to the overall effectiveness of the proposed framework.
@@ 3.1.2 Explainable Feature Extractor
Fig. 2.: Impact of sliding window size on the model performance. The optimal range (350–400) achieves a peak F1 score of 97%, balancing detection accuracy and temporal context.
@@
To validate the impact of the sliding window size on model performance, we conducted an experimental analysis using varying window sizes ranging from 0 to 1000. The results, depicted in Figure 3.1.2, indicate that a window size in the range of 350–400 achieves the optimal balance between detection performance and temporal context, yielding the highest F1 score of 97%. While larger windows resulted in marginal declines in performance due to excessive data aggregation, smaller windows struggled to capture meaningful temporal patterns, leading to reduced detection accuracy.
It is also important to highlight that the sliding window size has no impact on computational time or inference speed. This is because the sliding window operation simply involves adding the most recent packet and removing the oldest one from the window, making it a lightweight computation. As a result, the size of the sliding window does not affect the computational complexity or introduce latency during real-time detection. This ensures that the framework maintains its efficiency regardless of the selected window size while still benefiting from optimized detection performance.
@@ 3.1.5 Integrated Gradient Explainer
Fig. 4.: Illustration of structured prompt formats used in the zero-shot prompting. The left side shows the design and example of the flow-based prompt ($Q_{\text{flow}}$), where the predicted attack and contributing flow features are provided for explanation. The right side displays the structure and example of the payload-based prompt ($Q_{\text{payload}}$), where the LLM is asked to assess whether a given payload string is malicious or benign. For attacks that are payload-dependent, the final explanation generated by the LLM is derived by combining the responses from both $Q_{\text{flow}}$ and $Q_{\text{payload}}$, enabling a more comprehensive and context-aware interpretation.
@@ 3.1.6 Generative Explainer
While the LLM operates in a zero-shot setting, its ability to provide meaningful explanations is significantly enhanced by the newly developed temporal features and their contextual descriptions. These features enrich the prompt with domain-relevant information, guiding the LLM to interpret complex network behaviors more accurately. The structured nature of these features enables the LLM to produce precise, human-readable explanations and suggest remedial actions without requiring fine-tuning. However, we recognize that fine-tuning the LLM, especially using techniques like LoRA (Hu et al., 2022), could further refine its contextual understanding and tailor its responses to specific organizational needs. This enhancement is part of our future research agenda, aimed at improving both domain specific knowledge and organization specific remedial steps.
The development of well-structured prompts is a critical component of our zero-shot prompting approach, as it enables the LLM to generate accurate, context-aware explanations without requiring task-specific fine-tuning. The success of zero-shot inference is highly dependent on prompt quality, making it essential to design prompts that are both informative and aligned with the underlying data characteristics. The systematic process of constructing these prompts is described in Algorithm 3. To further clarify the structure and practical application of our prompt design, Fig. 3.1.5 presents visual examples of the prompt format, illustrating both the general structure and specific instances for payload-dependent and flow-dependent attacks. These examples help demonstrate how prompt formulation contributes to generating reliable and interpretable outputs from the LLM.
@@
These parts ($P_{\text{init}}$ , $P_{\text{part2}}$ , and $P_{\text{align}}$) are then combined to form a comprehensive query ($Q_{\text{flow}}$), which is then sent to the LLM to generate the flow-based response. If the GNN predicts a payload-specific attack, such as web-based or bruteforce attacks, additional processing is done for the payload data. In this processing, the payload importance vectors are normalized, and the top payloads are converted into a human-readable ASCII string. Consequently, a second query ($Q_{\text{payload}}$) is constructed to analyze the payload data. This query is prefixed with the phrase assigned to $P_{\text{payloadPrefix}}$: “Analyze whether this payload of network flow is malicious or not. Give reason concisely.” This query, along with the alignment section ($P_{\text{align}}$), is sent to the LLM to generate a corresponding response.
@@ 3.2 Dataset
To ensure a robust and comprehensive evaluation of the proposed XG-NID framework, we have employed three widely recognized benchmark datasets: CIC-IoT2023 (Neto et al., 2023), CIC-IDS2017 (Sharafaldin et al., 2018), and UNSW-NB15 (Moustafa and Slay, 2015). While the CIC-IoT2023 dataset is described in detail, including its preprocessing pipeline, we refrain from repeating similar preprocessing steps for CIC-IDS2017 and UNSW-NB15 to preserve readability and avoid redundancy. All three datasets were processed using a consistent pipeline to ensure fair comparison. Notably, since CIC-IDS2017 and UNSW-NB15 do not provide labeled PCAP files, we utilized the method introduced in Farrukh et al. (2022) to generate labeled flow-based data using the available flow metadata. This multi-dataset evaluation enhances the reliability and generalizability of our findings across diverse network traffic conditions and attack profiles.
3.3 CIC-IDS2017 Dataset
The CIC-IDS2017 dataset, developed by the Canadian Institute for Cybersecurity, captures a diverse range of modern network traffic scenarios, including both benign activities and sophisticated cyberattacks. The dataset simulates realistic traffic generated by 25 users over a period of five days, encompassing approximately 48.8 GB of network data. It is provided in both PCAP and CSV formats, with the latter comprising 80 flow-based features extracted using CICFlowMeter. The dataset includes over 2.8 million records across seven major attack categories: brute-force, DoS, DDoS, web-based attacks, infiltration, botnets, and port scanning. Several work in the literature has widely utilized this dataset for benchmarking intrusion detection research.
3.4 UNSW-NB15 Dataset
The UNSW-NB15 dataset, released by the Australian Centre for Cyber Security (ACCS), offers a rich collection of contemporary network traffic designed to emulate both legitimate behavior and a broad spectrum of attack vectors. Generated using the IXIA PerfectStorm tool within a controlled cyber range environment, the dataset captures over 31 hours of traffic distributed across 79 PCAP files, totaling more than 99 GB. It includes both raw packet data and flow-based representations generated using Argus and Bro-IDS, featuring over 45 extracted attributes. The dataset comprises approximately 2.5 million records and covers nine distinct attack types, including Fuzzers, Analysis, Backdoors, DoS, Exploits, Generic, Reconnaissance, Shellcode, and Worms. Although this dataset has been available for several years, it remains widely used by researchers as a benchmark for evaluating the performance of NIDS.
3.5 CIC-IoT2023 Dataset
@@ 3.5.1 Dataset Preprocessing
The second task involved addressing the significant class imbalance present in the CIC-IoT2023 dataset. Given the uneven distribution of classes, a combination of undersampling and oversampling techniques was employed to create a balanced dataset. Initially, 20% of the data samples were set aside to form the test set for each class. However, for classes with a large number of samples, the test set was capped at 4,000 randomly selected samples to avoid overrepresentation. In contrast, classes with fewer samples were retained as-is to preserve data integrity. The remaining samples were then allocated to the training set. To balance the training data, classes with more than 20,000 samples were undersampled by randomly selecting 20,000 samples, while classes with fewer than 20,000 samples were oversampled to reach 20,000 samples per class. This approach ensured a well-balanced distribution across all classes in both the training and test sets, as detailed in Table 4.
@@
Type Flows After Filter Test Set Train Set
@@
An important aspect of our preprocessing approach was the careful handling of attack subclasses. Since each main attack class in the CIC-IoT2023 dataset contains several subclasses, we ensured that the sampling process maintained proportional representation across these subclasses. This approach allowed us to preserve the diversity of attack types while achieving a more coherent and balanced dataset. Detailed information about the data preprocessing steps, including the implementation, can be found in the GitHub repository GNN4ID (Farrukh et al., 2024b).
@@ 4 Results and Analysis
This section presents a comprehensive evaluation of the proposed XG-NID framework, focusing on its detection performance, explainability, and computational efficiency in real-time network intrusion detection scenarios. The analysis is structured into two primary subsections: Performance Analysis and Explainability Analysis.
The Performance Analysis subsection provides a thorough comparison of the proposed framework against baseline models and state-of-the-art techniques, demonstrating its superiority and robustness in detecting network intrusions by effectively fusing packet-level and flow-level information. Additionally, this subsection includes a focused Scalability and Computational Overhead analysis, which assesses the framework’s feasibility for real-time intrusion detection. Key factors such as inference time and the influence of computational hardware on the overall security pipeline are discussed to highlight the system’s operational efficiency. Conversely, the Explainability Analysis subsection explores how our framework leverages contextual information to produce explainable and actionable insights, thereby enhancing the interpretability of the model’s predictions and facilitating informed decision-making by security operators.
All evaluations presented in this section are conducted in the context of multi-class classification. To ensure fairness, reliability, and reproducibility, all experiments were performed under consistent experimental conditions, with uniform data preprocessing and distribution strategies applied across all models. The approaches included for comparison were implemented by closely adhering to the methodologies outlined in their respective original papers, thereby preserving the integrity of their evaluation protocols while aligning them within a unified experimental framework.
@@ 4.1 Performance Analysis
Conducting a direct, apple-to-apple comparison of the proposed framework presents certain challenges, as, to the best of the authors’ knowledge, this is the first work to fuse dual modalities—packet-level and flow-level information of network traffic—within a HGNN. Consequently, comparing our framework with existing approaches that solely utilize either flow-level or packet-level information may not yield an accurate assessment of its overall effectiveness.
Furthermore, comparing our proposed HGNN with existing graph-based NIDS introduces inherent challenges. Most prior works employing Graph Neural Networks are designed for edge-level or node-level classification, often relying on historical traffic analysis rather than enabling real-time inference, as discussed in Section 2. In contrast, our framework is explicitly developed for graph-level classification with real-time detection capabilities. As a result, direct comparisons with node- or edge-level models are not directly applicable due to fundamental differences in problem formulation and system objectives.
Given that our framework is tailored for real-time inference while effectively fusing dual modalities to improve the detection of both flow-level and packet-level attacks, a tailored evaluation strategy is essential. To ensure a comprehensive assessment of our approach, we divided the performance analysis into two distinct parts.
@@
Finally, we present a concise analysis of the computational efficiency and scalability of the proposed framework. This assessment covers key factors such as inference time, the impact of resource utilization on overall framework performance, and the additional overhead introduced by integrating the LLM for generating explainable and actionable insights. By examining these aspects, we aim to showcase the framework’s capability to support real-time intrusion detection while ensuring optimal operational efficiency.
@@ 4.1.1 Baseline Comparisons
In evaluating the baseline models, our goal is to present a comprehensive comparison that highlights the advantages of the proposed framework over approaches that utilize only flow-level or packet-level information. For clarity and conciseness, baseline comparisons were conducted exclusively on the CIC-IoT2023 dataset. To further enhance the analysis, we examined the performance of these baseline models against specific categories of attacks—specifically, payload-dependent attacks (e.g., brute-force and web-based attacks) using flow-level features, and payload-independent attacks using packet-level features. This breakdown provides deeper insights into the strengths and limitations of single-modality approaches in contrast to our unified dual-modality framework.
@@
Fig. 5.: Performance comparison between the proposed framework and several state-of-the-art models that incorporate dual modalities of network traffic, evaluated across three benchmark datasets (CIC-IoT2023, CIC-IDS2017, and UNSW-NB15). The proposed framework consistently outperforms existing approaches, achieving up to 97% F1-score—a widely accepted metric for assessing multiclass classification performance in intrusion detection systems.
@@
McPID (Hojjatinia et al., 2023) 0.71 0.65 0.68 0.60
Parallel ViT (Zhang, 2024) 0.82 0.78 0.80 0.68
BI-TBL (Wanshun et al., 2023) 0.78 0.74 0.71 0.65
@@ 4.1.2 Dual-Modality State-of-the-Art
As illustrated in Fig. 4.1.1, our proposed framework outperforms several state-of-the-art methods in terms of detection performance. To ensure a comprehensive and robust evaluation, we assessed our approach across all three benchmark datasets, providing a broader perspective on its generalizability. In Fig. 4.1.1, we report results using the F1-score, which offers a more balanced and representative metric for multiclass classification tasks. The superior performance of our model can be attributed to its inherent fusion of packet-level and flow-level information, enabling more precise and efficient classification. This integration not only enhances detection accuracy but also provides more actionable insights, establishing our framework as a cutting-edge solution in the domain of network intrusion detection.
@@
Table 7: Overview of Component-Wise Performance and Influencing Factors in the XG-NID Framework.
Component Influencing Factors Model Measured Performance
Network Packet Capture - Network Bandwidth Python (Scapy) 26 packets/second
- NIC and CPU Performance
- Buffer Sizes
Real-time Detection - GPU Capability HGNN 6.563 ms/sample
- Preprocessing Pipeline
- Model Complexity
LLM Inference - Model Size Llama 3-8B $\sim$64 tokens/second ($\sim$6s per 390-token explanation)
- GPU Capability
- Prompt Complexity DeepSeek-V2 Lite $\sim$20 tokens/second ($\sim$19.5s per 390-token explanation)
- Model Quantization
4.1.3 Scalability and Computational Overhead
Scalability and computational overhead are critical considerations in evaluating the real-time feasibility of the proposed XG-NID framework, especially given the integration of a LLM for generating explainable and actionable insights. While the LLM enhances interpretability, its inclusion raises concerns about real-time performance, as the speed of network packet capture significantly exceeds typical LLM response times. However, it is important to note that both packet capture rates and LLM inference speeds are influenced by several factors. As summarized in Table 7, multiple factors influence both the network packet capture rates and LLM inference speeds. For packet capture using Python’s Scapy, network bandwidth, NIC and CPU performance (Marinos et al., 2014), and buffer sizes are the main influencing factors. Similarly, the inference speed of the LLM (Llama 3-8B and DeepSeek-V2 Lite) is influenced by several factors, including model size, GPU capabilities, prompt complexity, and model quantization.
Despite these differences, the LLM in XG-NID serves a supplementary role, providing post-detection explanations without affecting the core real-time detection pipeline, which is handled exclusively by the HGNN. Additionally, the LLM is only triggered for samples flagged as attacks, ensuring it does not continuously burden the system.
To quantify computational efficiency, we measured the HGNN’s inference latency, which averaged approximately 6.563 milliseconds per sample, as reported in Table 7. For the LLMs, the average explanation spanned around 390 tokens, with an observed inference speed of 64 tokens per second for LLaMA-3 8B and 19 tokens per second for DeepSeek-V2 Lite. It is important to note that both models were utilized in Q4_0 quantized format, following the default configuration of the Ollama framework (Ollama, 2024), which significantly reduces memory consumption and allows execution on a consumer-grade GPUs (RTX 3070 GPU, 8GB). Consequently, while attack detection occurs within milliseconds, a detailed explanation is provided to analysts within approximately 6 seconds using LLaMA-3 8B—though this latency can be reduced with enhanced computational resources.
Additionally, packet capture experiments using the same hardware setup recorded an average of 26 packets per second. This provides a practical baseline to evaluate real-time capabilities, highlighting the balance between detection speed and interpretability. Furthermore, the disparity between packet capture rates and LLM inference speeds emphasizes the importance of optimized system design.
Finally, it is essential to acknowledge that operational speeds of practical Network-Based Intrusion Detection Systems frequently remain below actual network speeds (Lai et al., 2004) due to deep packet inspection’s computational intensity. However, XG-NID’s real-time performance can substantially improve through high-end hardware deployment, parallel processing techniques, and further algorithmic optimizations, ensuring scalability while maintaining operational efficiency.